OpenStudy (anonymous):

There are 4 blue marbles and 3 red marbles in a bag. You randomly pick one marble, keep it, and then pick a second marble. What is the probability that you will pick two blue marbles?

OpenStudy (anonymous):

2/7 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

8/21 ? @satellite73

OpenStudy (anonymous):

oh there you are

OpenStudy (anonymous):

what is the probability that the first one selected is blue?

OpenStudy (anonymous):

4/7 ? @satellite73

OpenStudy (anonymous):

yes now that marble is in your pocket what is the probability that the second marble selected is blue now that one blue marble is gone?

OpenStudy (anonymous):

oh heavens, i see you had the right answer to begin with!!

OpenStudy (anonymous):

the same right because everytime i take one out i put it back in there ? @satellite73

OpenStudy (anonymous):

and me, like a dope, said "no" so sorry

OpenStudy (anonymous):

no now it is \(\frac{3}{6}=\frac{1}{2}\) because one blue marble is gone and there are now 3 blue and 6 total so you were right \[\frac{4}{7}\times \frac{1}{2}=\frac{2}{7}\]

OpenStudy (anonymous):

you were right, sorry to waste your time i misread blue for red you had it on the first try

OpenStudy (anonymous):

its cool so its 2/7

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

lol for what it is worth, you are welcome

OpenStudy (anonymous):